Retained Earnings Question

You have been given the following information for PattyCake’s Athletic Wear Corp. for the year 2018:

  1. Net sales = $38,700,000.

  2. Cost of goods sold = $22,180,000.

  3. Other operating expenses = $6,000,000.

  4. Addition to retained earnings = $1,206,500.

  5. Dividends paid to preferred and common stockholders = $1,933,000.

  6. Interest expense = $1,830,000.

  7. The firm’s tax rate is 30 percent.

In 2019:

  1. Net sales are expected to increase by $9.70 million.

  2. Cost of goods sold is expected to be 60 percent of net sales.

  3. Depreciation and other operating expenses are expected to be the same as in 2018.

  4. Interest expense is expected to be $2,105,000.

  5. The tax rate is expected to be 30 percent of EBT.

  6. Dividends paid to preferred and common stockholders will not change.

Calculate the addition to retained earnings expected in 2019.

Thank you for your help!
Answer #1
working Amount in 2019
sales 38700000+9700000 48400000
cost of goods sold 48400000*60% 29040000
Gross profit 19360000
Depreciation 4205000
Other operating expense 6000000
Income before interest and tax 9155000
less:Interest expense 2105000
Income before tax 7050000
less:Tax expense 7050000*30% 2115000
Net income 4935000
less:Preferred and equity dividend 1933000
Addition to retained earning 3002000

working:

Net income in 2018 =Preferred and equity dividend +addition to retained earning

       = 1933000+1206500

            = 3139500

Income before tax = Net income /[1-tax rate]

              = 3139500/ [1-.30]

             = 4485000

Income before interest and tax = 4485000+1830000 = 6315000

Income before interest and tax =sales -cost of sales -depreciation -operating expense

6315000 = 38700000-22180000-Depreciation-6000000

6315000 = 10520000-depreciation

Depreciation = 10520000-6315000

               = 4205000